OpenStudy (anonymous):

a2 + b2 = 5ab find the value of a3/b3 + b3/a3

OpenStudy (anonymous):

use \[(x^3+y^3)=(x+y)(x^2-xy+y^2)\]

OpenStudy (anonymous):

a/b=x b/a=y a/b+b/a=5

OpenStudy (anonymous):

x+y=5